OpenStudy (anonymous):

Find the indefinite integral of ((1/2t)-(2)^(1/2)e^t)dt

OpenStudy (turingtest):

\[\large\int\frac12t-2^{\frac12e^t}dt\]This?

OpenStudy (anonymous):

No, the first t is in bottom with the 2, and the second 2 is under a radical by itself and e^t is another variable not an entire exponent

OpenStudy (anonymous):

\[\Large \int \frac{1}{2t}dt- 2^{1/2} \int e^t dt\] If that's not what you mean I suggest you, trying to \( \LaTeX\) it out yourself, so we all talk about the same problem.

OpenStudy (anonymous):

It's almost right just include e^t after the square root of 2 and take off dt till the end

OpenStudy (anonymous):

\[\int\limits_{-\infty}^{\infty}[\frac{ 1 }{ 2t }-\sqrt{2}e^t]dt\]

OpenStudy (anonymous):

is that your question?

OpenStudy (anonymous):

@Spacelimbus is right you need to tell him the right question so he can help you to provide you with the right solution.

OpenStudy (anonymous):

\[\int\limits[\frac{ 1 }{ 2t }-\sqrt{2}e^t]dt\]

OpenStudy (anonymous):

sorry this is an indefinite integral i think you are asking

OpenStudy (anonymous):

if this is the right question then the answer would be

OpenStudy (anonymous):

\[=\frac{ 1 }{ 2 }\ln(t)-2^\frac{ 1 }{2 } e^t+c\]

OpenStudy (anonymous):

otherwise help us to help you by providing the right question

OpenStudy (anonymous):

Yes that's the right question! But how you solve it?

OpenStudy (anonymous):

Forget it, I got it, thank you